chocolateluva77:

can smo reply asap? without solving the system, determine whether there are 0, 1 or an infinite number of solutions {-4x + 2y = -14 y = 2x + 8

umm:

Hey there, @chocolateluva77! Is there a way you may provide a screenshot?

chocolateluva77:

sure

chocolateluva77:

chocolateluva77:

question 5

chocolateluva77:

@sslpj do yk how to do it?

TwilightQueenBee:

Take all the terms to the LHS and name the constants in equation 1 as a1, b1, c1 and that in second equation as a2, b2, c2. If a1/a2 not equal to b1/b2 --- One unique solution If a1/a2 = b1/b2 = c1/c2 --- Infinite solutions If a1/a2 = b1/b2 not equal to c1/c2 --- No solution

chocolateluva77:

what does LHS mean?

TwilightQueenBee:

Left hand side Left side of your equation.

chocolateluva77:

@TwilightQueenBee can u do it on a paper with the steps if u don't mind n say if there's 0, 1 or a infinite number of solutions and post it on here?

chocolateluva77:

@twilightqueenbee can u help?